# Why is Grain Growth Not Curvature Flow?

### v1

@@ IV Methods
$\begin{split}F=\int_{\Omega}{\rm d}\Omega\bigg{[}&\sum_{i}^{N}\sum_{j}^{N}\frac{4\gamma_{ij}}{\epsilon}\bigg{(}\eta_{i}\eta_{j}-\frac{\epsilon^{2}}{\pi^{2}}\nabla\eta_{i}\cdot\nabla\eta_{j}\bigg{)}\\ &+\sum_{i}^{N}\eta_{i}\psi_{i}+\lambda\bigg{(}\sum_{i}^{N}\eta_{i}-1\bigg{)}\bigg{]},\end{split}$ (3)
with $\gamma_{ij}$ the interface/grain-boundary energy, $\epsilon$ a parameter controlling the extension of the diffuse interface between grains, $\psi_{i}$ the bulk free energy, and $\lambda$ a Lagrange multiplier. The double sum approximates the interface energy, and the last term is a constraint, imposing that in every point in $\Omega$, the sum of all the phase fields is 1 (i.e., enforcing complete filling of the domain with grains). Leveraging the definition of the free energy (3), the motion of GBs dictated by Eq. (2) can be approximated by the following evolution law of $\eta_{i}$:
$\dot{\eta}_{i}=\sum_{j}^{N}M_{ij}\left(\frac{\delta F}{\delta\eta_{i}}-\frac{\delta F}{\delta\eta_{j}}+\mathbf{f}_{\rm elastic}\right).$ (4)
@@
$\frac{\delta F}{\delta\eta_{i}}-\frac{\delta F}{\delta\eta_{j}}=\mathbf{f}_{\rm curvature}+\mathbf{f}_{\rm chemical},$ (5)
$\mathbf{f}_{\rm curvature}=\Gamma_{ij}\left(\eta_{j}\nabla^{2}\eta_{i}-\eta_{i}\nabla^{2}\eta_{j}-\frac{\pi^{2}}{2\epsilon^{2}}(\eta_{j}-\eta_{i})\right),$ (6)
$\mathbf{f}_{\rm chemical}=\frac{\pi}{\epsilon}\sqrt{\eta_{i}\eta_{j}}(\psi_{j}-\psi_{i}),$ (7)
$\mathbf{f}_{\rm elastic}=\boldsymbol{\tau}\cdot\boldsymbol{\beta}_{ij}|\nabla\eta_{j}|(\hat{\mathbf{n}}(\eta_{i})\cdot\hat{\mathbf{n}}(\eta_{j}))$ (8)
and $\hat{\mathbf{n}}(\eta)=-\nabla\eta/|\nabla\eta|$, accounting for the elastic interactions with external stress or disconnection self-stress via $\boldsymbol{\tau}$, obtained computing the corresponding sharp-interface quantity $\boldsymbol{\tau}$ and extending it off the interface like in level-set approaches [24]. Numerical simulations are performed exploiting a simple finite-difference spatial discretization of Eq. (4) and semi-implicit time integration. Further details concerning model parametrization and additional numerical simulations are reported in the SI.
Acknowledgements.
@@
Why is Grain Growth Not Curvature Flow?
@@
The theoretical framework considered in this work builds on a continuum model for disconnection-mediated grain boundary motion that respects the underlying bicrystallography [1, 2, 3]. We briefly summarize the continuum model as follows. A grain boundary (GB) is modeled as a continuous curve $\Sigma$, described by $\mathbf{x}(s)$ in the $\hat{\mathbf{e}}_{1}-\hat{\mathbf{e}}_{2}$ two-dimensional plane with $s$ a parameter. Low energy orientation for $\Sigma$, namely low energy GBs, exist, which are considered reference orientations $R(k)$ with $k=0,\dots,N-1$ and $N$ the number of considered references. For instant, the [110]-tilt GBs in FCC crystals possess two reference interfaces ($N=2$); the [100]-tilt GBs exhibit four references ($N=4$); the [111]-tilt GBs have six references ($N=6$). An arbitrarily oriented $\Sigma$ may be represented as the nearest reference interfaces with a superimposed distribution of disconnections with step character $h^{(k)}$ and Burgers vector $\mathbf{b}^{(k)}$ [4]. The abundance of disconnections is quantified by the respective densities $\rho^{(k)}$ and follows from geometrical considerations, namely how many disconnections per unit length are needed to obtain the actual orientation of $\Sigma$ from the closest reference orientations. They can be thus related to the unit tangent vector $\mathbf{l}$ of $\Sigma$,
@@
with $\hat{\mathbf{e}}^{(k)}$ the orientations of $k$-th reference, with $k$ and $k+1$ labelling the references with the closest orientations to $\Sigma$. The equation of motion for the latter can be derived from the corresponding definition of the interface energy assuming gliding of disconnections only and overdamped dynamics. It results
@@
with $\hat{\mathbf{n}}=\mathbf{l}^{\prime}(s)$ the vector normal to $\Sigma$, $\mathbf{M}$ the mobility tensor, $\Gamma=\gamma(\phi)+\gamma^{\prime\prime}(\phi)$ the interface stiffness with $\gamma(\phi)$ the energy density of the interface and $\phi$ the inclination angle, $\kappa=\nabla_{s}\hat{\mathbf{n}}$ the interface curvature, $\boldsymbol{\tau}=(\tau^{(k)},\tau^{(k+1)})$ the resolved shear stress vector at $\mathbf{x}$, $\boldsymbol{\beta}$ the shear coupling factors, $\Psi$ the jump in the chemical potential, and $\hat{\mathbf{n}}=$ the vector normal to $\Sigma$. Parameters follow from the corresponding values for references $R(k)$ and $R(k+1)$: $\mathbf{M}\equiv(M^{(k)},M^{(k+1)})\mathbf{I}$, $\boldsymbol{\beta}\equiv(\beta^{(k)},\beta^{(k+1)})\mathbf{I}$, $\gamma(\phi)\equiv\gamma^{(k)}|\sin(\phi-\phi^{(k)})|/|\sin(\Delta\phi^{(k)})|+\gamma^{(k+1)}|\sin(\phi^{(k+1)}-\phi)|/\sin(\Delta\phi^{(k)})$, with $\Delta\phi^{(k)}=\phi^{(k+1)}-\phi^{(k)}$ (and note that $\phi^{(k)}<\phi<\phi^{(k+1)}$ by construction).
The three terms in Eq. (S-2) account for capillarity-driven GB motion (first term), coupling with (shear) stress (second term), and bulk-energy differences across GBs (third term). The shear stress includes both the effect of applied external fields and the self-stress generated by the dislocation character of all the disconnections on GBs in the system. Explicitly, $\tau^{(k)}=(\sigma_{22}-\sigma_{11})\sin(2\phi^{(k)})/2+\sigma_{12}\cos(2\phi^{(k)})$ with $\sigma_{ij}=\sigma_{ij}^{\rm ext}+\sigma_{ij}^{\rm self}$ ($\{i,j\}\in\{1,2\}$). $\sigma_{ij}^{\rm self}$ can be approximated as the stress field generated by a density of dislocations with Burgers vector $\mathbf{b}^{(k)}$ by along the entire interface given by
$\sigma_{ij}^{\rm self}(s)=K\sum_{k}^{N}\int_{\Sigma}|\mathbf{b}^{(k)}|\rho^{(k)}(s_{0})\widetilde{\sigma}_{ij}(\mathbf{x}(s)-\mathbf{x}(s_{0})){\rm d}s_{0}.$ (S-3)
with $K=\mu/[2\pi(1-\nu)]$.
The continuum model described so far applies to bicrystal configurations with one GB ($\Sigma$). However, it can be considered for microstructures featuring a network of grain boundaries, by defining parameters and geometrical properties accordingly. This is conveniently realized in the implicit description of GBs below, allowing us also to cope easily with complex morphological evolution (typical of grain growth).
@@
$\dot{\eta}_{i}=\sum_{j}^{N}M_{ij}\left(\frac{\delta F}{\delta\eta_{i}}-\frac{\delta F}{\delta\eta_{j}}+\mathbf{f}_{\rm elastic}\right).$ (S-5)
@@
$\frac{\delta F}{\delta\eta_{i}}-\frac{\delta F}{\delta\eta_{j}}=\mathbf{f}_{\rm curvature}+\mathbf{f}_{\rm chemical},$ (S-6)
$\mathbf{f}_{\rm curvature}=\Gamma_{ij}\left(\eta_{j}\nabla^{2}\eta_{i}-\eta_{i}\nabla^{2}\eta_{j}-\frac{\pi^{2}}{2\epsilon^{2}}(\eta_{j}-\eta_{i})\right),$ (S-7)
$\mathbf{f}_{\rm chemical}=\frac{\pi}{\epsilon}\sqrt{\eta_{i}\eta_{j}}(\psi_{j}-\psi_{i}),$ (S-8)
$\mathbf{f}_{\rm elastic}=\boldsymbol{\tau}\cdot\boldsymbol{\beta}_{ij}|\nabla\eta_{j}|(\hat{\mathbf{n}}(\eta_{i})\cdot\hat{\mathbf{n}}(\eta_{j}))$ (S-9)
@@
Numerical simulations are performed exploiting a finite difference spatial discretization of (S-5) and semi-implicit time integration.
The space is discretized into rectangles with edges ($\mathrm{d}x_{1}$ and $\mathrm{d}x_{2}$). In all current PF simulations, $\mathrm{d}x_{1}=\mathrm{d}x_{2}=\epsilon/5$. The discretised gradient and Laplacian operators are
@@
where $l$ and $m$ are the index of the element along the $\mathbf{e}_{1}$ and $\mathbf{e}_{2}$ directions.
All the simulations are performed under periodic boundary condition in both $\mathbf{e}_{1}$ and $\mathbf{e}_{2}$ directions.
@@
We generate 30 samples for the PF simulations with different initial configuration. The initial configuration for grain growth simulations is constructed with 1000 circle nuclei with different orientations within a background “substrate”. A chemical potential difference between the nuclei and the background “substrate” is applied so that the nuclei can rapidly consume the background “substrate” and densify into a polycrystal microstructure. We assume that the 2D polycrystals are [110]-textured in face-centered cubic metals; i.e., two sets of closely-packed and low-energy GB planes exist and are perpendicular to each other in the CSL lattices of [110]-tilt GBs. If we set the orientation $[001]\,||\,\mathbf{e}_{x}$, $[1\bar{1}2]\,||\,\mathbf{e}_{y}$ and $[110]\,||\,\mathbf{e}_{z}$ as the reference, according to crystal symmetry, the orientation angles of all grains should be in range of [0, $\pi$).
For the cases of mean curvature flow and those considering GB shear coupling (shown in Figs. 1-3 and 4f-g in main text), the GB energy ($\gamma$) and mobility ($M$) of all GBs are set as $\gamma=1$ and $M=4\pi^{2}/8\epsilon\approx 10$, where $\epsilon=5\mathrm{d}x$ is the interface thickness in the PF framework. Shear coupling factor of the reference interfaces of a certain arbitrarily-curved GB is determined by their misorientation angle ($\theta$) according to [4]
@@
We note that shear coupling factors measured in polycrystals tend to be significantly smaller than those found in bicrystals [9, 10, 11, 20]. We account for this difference by introducing a shear coupling scaling factor, $S=1/30$. While the value of $S$ is chosen arbitrarily, it is of the correct order of magnitude compared with experimental observations. Nonetheless, the shear coupling factor expression, Eq. (S-12), retains the dependence of $\beta$ on bicrystallography consistent with theory, MD simulations and experiments. The source of the smaller shear coupling in polycrystals compared with bicrystals is associated with the constraints imposed by other grains (i.e., triple junctions) in a polycrystal that activate additional disconnection modes [12, 13, 14].
@@
As for the [001]-textured idealized microstructure, the initial configuration is also composed of square grains A and B and octagon grains C and D with grains A and B sharing the same orientation –i.e., $[100]\,||\,\mathbf{e}_{x}$, $[010]\,||\,\mathbf{e}_{y}$ and $[001]\,||\,\mathbf{e}_{z}$. Grain C has a misorientation angle -22.6 away from grains A and B. Grain D has a misorientation angle 28.1 away from grains A and B. In this sense, the GBs between grains A and C, B and C are $\Sigma 13[100]$ asymmetric tilt GBs, whilst the GBs between grains A and D, B and D are $\Sigma 17[100]$ asymmetric tilt GBs. The shear coupling factors are also assigned to the inidividual GBs based on their misorientations  [11].
@@
In the Figs.4 h,i of the main text, we introduce the misorientation-dependent GB energy. The formulation of GB energy follows the generalized GB energy functional proposed by Bulatov, Reed and Kumar [15].
@@
Figure S-1: Results on anisotropic grain growth considering anisotropic GB energy and triple junction drag effect. (a) Mean grain size $\langle R^{2}\rangle$ vs. time. (b) Mean grain size $\langle R\rangle$ vs. time. Scatter plots of grain area change $\delta R^{2}/\delta t$ vs. number of neighbours $N$ of grains during grain growth considering (c) anisotropic GB energy and (d) additional smaller triple junction mobility.
@@
The triple junction (TJ) drag effect is introduced by a lower PF kinetic coefficient in the TJ region (i.e., TJ mobility) than that in the GB region (i.e., GB mobility) [16].
How to find the TJ region? If the grid locates inside the grain $i$ or inside the grain $j$ or at the GB between them, $\eta_{i}(l,m)+\eta_{j}(l,m)=1$. The TJ region locates at the grid ($l,m$) where $\eta_{i}(l,m)+\eta_{j}(l,m)<1$. Then, the mobility $M_{ij}$ at grid ($l,m$) in Eq. (S-5) should be
$M_{ij}=M_{ij}^{GB}(M_{ij}^{TJ}+\exp(-500(\eta_{i}+\eta_{j}-1)^{2})(1-M_{ij}^{TJ})).$ (S-14)
It is clear that if $\eta_{i}+\eta_{j}=1$ (at GBs), $M_{ij}=M_{ij}^{GB}$; if if $\eta_{i}+\eta_{j}\ll 1$ (at TJs), $M_{ij}=M_{ij}^{TJ}$. Here, in this work, we choose $M^{TJ}=0.1M^{GB}$.
Figure S-2: Illustration of the GB replacement happening in one typical grain. The GBs are colored based on their GB energy.
@@
In the main text, we mention that we conduct PF simulations on the grain growth considering no GB shear coupling but with anisotropic GB energy and triple junction (TJ) drag effect.
[…]
@@
Figures S-1c and d show the scatter plots of grain size changes vs. number of neighbours. The thick gray lines refer to the ideal von Neumann relation. It is clear that the data distribution deviates away from the ideal von Neumann relation. Although the deviation in grain size changes vs. number of neighbours relation, the distribution of data (and the shapes of the up limit and lower limit) is still largely different from the data distribution in experiments. However, clear linear distributions are found and the only difference is the reduction of the slope of the linear distributions.
@@
In previous studies [17, 18], the reason for the uncorrelatedness between GB velocity and mean curvature in polycrystalline Ni and $\alpha$-Fe is the occurance of GB replacement due to GB energy anisotropy. We demonstrate that GB velocity is still correlated with mean curvature even considering GB energy anisotropy. Here, we show that GB replacement can also be found in our PF simulations, see Fig. S-2. High energy boundaries 1 and 3 disappear during evolution, whilst the low energy boundaries 2 and 7 increase their lengths. The grain size increases less than 1% during this procedure. This phenomenon can be described as ”GB replacement” according to the definition in [17, 18].

### v3

@@ III Methods
$\begin{split}F=\int_{\Omega}{\rm d}\Omega\bigg{[}&\sum_{i}^{N}\sum_{j}^{N}\frac{4\gamma_{ij}}{\epsilon}\bigg{(}\eta_{i}\eta_{j}-\frac{\epsilon^{2}}{\pi^{2}}\nabla\eta_{i}\cdot\nabla\eta_{j}\bigg{)}\\ +&\sum_{i}^{N}\eta_{i}\psi_{i}+\lambda\bigg{(}\sum_{i}^{N}\eta_{i}-1\bigg{)}\bigg{]},\end{split}$ (3)
with $\gamma_{ij}$ the interface/grain-boundary energy density, $\epsilon$ a parameter controlling the extension of the diffuse interface between grains, $\psi_{i}$ the bulk free energy, and $\lambda$ a Lagrange multiplier. The first term approximates the interface energy, and the last term is a constraint, imposing that in every point in $\Omega$, the sum of all the phase fields is 1 (i.e., enforcing complete filling of the domain with grains). Leveraging the definition of the free energy Eq. (S-4), the motion of GBs dictated by Eq. (2) can be approximated by the following evolution law of $\eta_{i}$:
$\dot{\eta}_{i}=\sum_{j}^{N}M_{ij}\left(\frac{\delta F}{\delta\eta_{j}}-\frac{\delta F}{\delta\eta_{i}}+f_{\rm elastic}\right).$ (4)
@@
$\frac{\delta F}{\delta\eta_{j}}-\frac{\delta F}{\delta\eta_{i}}=f_{\rm curvature}+f_{\rm chemical},$ (5)
$f_{\rm curvature}=\Gamma_{ij}\left(\eta_{j}\nabla^{2}\eta_{i}-\eta_{i}\nabla^{2}\eta_{j}-\frac{\pi^{2}}{2\epsilon^{2}}(\eta_{j}-\eta_{i})\right),$ (6)
$f_{\rm chemical}=\frac{\pi}{\epsilon}\sqrt{\eta_{i}\eta_{j}}(\psi_{j}-\psi_{i}),$ (7)
$f_{\rm elastic}=\boldsymbol{\tau}\cdot\boldsymbol{\beta}_{ij}|\nabla\eta_{j}|(\hat{\mathbf{n}}(\eta_{i})\cdot\hat{\mathbf{n}}(\eta_{j}))$ (8)
and $\hat{\mathbf{n}}(\eta)=-\nabla\eta/|\nabla\eta|$, accounting for the elastic interactions with external stress or disconnection self-stress via $\boldsymbol{\tau}$. This is obtained by computing the corresponding sharp-interface quantity and extending it off the interface like in Level-set approaches [27]. The continuum model of interface migration mediated by disconnection flow, Eq. (2), and thus its PF formulation, follows self-consistently from the microscopic picture of arbitrarily curved GBs with disconnections [26, 27, 29]. The density of disconnections is dictated by the geometry of the interface on which they flow, as well as dynamical/energetic effects (in a multiple-disconnection-mode formulation); see SI. In the original formulation [26, 27, 29], parameters such as GB mobility, energy, and shear coupling factors, are expressed as a function of those for reference orientations (which define the expected disconnections). In this work, we set directly the shear coupling factor of the reference of arbitrarily curved GBs as a function of their misorientation [23]; the corresponding full expression is reported in the SI. This description is shown to be fully in agreement with the microscopic picture of GB migration in the Discussion section. Numerical simulations are performed exploiting a simple finite-difference spatial discretization of Eq. (S-5) and semi-implicit time integration. Further details concerning model parametrization and additional numerical simulations are reported in the SI.
Acknowledgements
@@
Why Grain Growth is Not Curvature Flow
@@
The theoretical framework considered in this work builds on a continuum model for disconnection-mediated grain boundary (GB) motion that respects the underlying bicrystallography [1, 2, 28]. We briefly summarize the continuum model as follows. A GB is modeled as a continuous curve $\Sigma$, described by $\mathbf{x}(s)$ in the $\hat{\mathbf{e}}_{1}-\hat{\mathbf{e}}_{2}$ two-dimensional plane with $s$ a parameter. Low energy orientation for $\Sigma$, namely low energy GBs, exist, which are considered reference orientations $R(k)$ with $k=0,\dots,N-1$ and $N$ the number of considered references. For instance, the [110]-tilt GBs in FCC crystals possess two reference interfaces ($N=2$), the [100]-tilt GBs exhibit four references ($N=4$), and the [111]-tilt GBs have six references ($N=6$). An arbitrarily oriented $\Sigma$ may be represented as the nearest reference interfaces with a superimposed distribution of disconnections with step character $h^{(k)}$ and Burgers vector $\mathbf{b}^{(k)}$ [4]. The abundance of disconnections is quantified by the respective densities $\rho^{(k)}$ and follows from geometrical considerations, namely how many disconnections per unit length are needed to obtain the actual orientation of $\Sigma$ from the closest reference orientations. They can be thus related to the unit tangent vector $\mathbf{l}$ of $\Sigma$,
@@
with $\hat{\mathbf{e}}^{(k)}$ the orientations of $k$-th reference, with $k$ and $k+1$ labelling the references with the closest orientations to $\Sigma$. The equation of motion for the latter can be derived from the corresponding definition of the interface energy, assuming gliding of disconnections only and overdamped dynamics. It results
@@
with $\hat{\mathbf{n}}=\mathbf{l}^{\prime}(s)$ the vector normal to $\Sigma$, $\mathbf{M}$ the mobility tensor, $\Gamma=\gamma(\phi)+\gamma^{\prime\prime}(\phi)$ the interface stiffness with $\gamma(\phi)$ the energy density of the interface and $\phi$ the inclination angle, $\kappa=\nabla_{s}\hat{\mathbf{n}}$ the interface curvature, $\boldsymbol{\tau}=(\tau^{(k)},-\tau^{(k+1)})^{\mathsf{T}}$ the resolved shear stress vector at $\mathbf{x}$, $\boldsymbol{\beta}$ the shear coupling factors, $\psi$ the jump in the chemical potential, and $\hat{\mathbf{n}}$ the vector normal to $\Sigma$. Parameters follow from the corresponding values for references R$(k)$ and R$(k+1)$: $\mathbf{M}\equiv(M^{(k)},M^{(k+1)})^{\mathsf{T}}$, $\boldsymbol{\beta}\equiv(-\beta^{(k)},\beta^{(k+1)})^{\mathsf{T}}$, $\gamma(\phi)\equiv\gamma^{(k)}|\sin(\phi-\phi^{(k)})|/|\sin(\Delta\phi^{(k)})|+\gamma^{(k+1)}|\sin(\phi^{(k+1)}-\phi)|/\sin(\Delta\phi^{(k)})$, with $\Delta\phi^{(k)}=\phi^{(k+1)}-\phi^{(k)}$ (and note that $\phi^{(k)}<\phi<\phi^{(k+1)}$ by construction).
The three terms in Eq. (S-2) account for capillarity-driven GB motion (first term), coupling with (shear) stress (second term), and bulk-energy differences across GBs (third term). The shear stress includes both the effect of applied external fields and the self-stress generated by the dislocation character of all the disconnections on GBs in the system. Explicitly, $\tau^{(k)}=(\sigma_{22}-\sigma_{11})\sin(2\phi^{(k)})/2+\sigma_{12}\cos(2\phi^{(k)})$ with $\sigma_{ij}=\sigma_{ij}^{\rm ext}+\sigma_{ij}^{\rm self}$ ($\{i,j\}\in\{1,2\}$). $\sigma_{ij}^{\rm self}$ can be approximated as the stress field generated by a density of dislocations with Burgers vector $\mathbf{b}^{(k)}$ along the entire interface given by
$\sigma_{ij}^{\rm self}(s)$ $=K\sum_{k}^{N}\int_{\Sigma}|\mathbf{b}^{(k)}|\rho^{(k)}(s_{0})\widetilde{\sigma}_{ij}(\mathbf{x}(s)-\mathbf{x}(s_{0})){\rm d}s_{0}$ (S-3)
$=K\sum_{k}^{N}\int_{\Sigma}\beta^{(k)}\big{[}h^{(k)}\rho^{(k)}(s_{0})\big{]}\widetilde{\sigma}_{ij}(\mathbf{x}(s)-\mathbf{x}(s_{0})){\rm d}s_{0}$
with $K=\mu/[2\pi(1-\nu)]$ and $\widetilde{\sigma}_{ij}$ are the stress field components of a single disconnection (assuming that all disconnections only have edge dislocation character) [2]. Notice that we do not directly model disconnections explicitly [2, 5, 3]. The term $h^{(k)}\rho^{(k)}$ is determined by local tangent in Eq. (S-1) instead of real disconnection step heights, and the shear coupling factor effectively accounts for the Burgers vectors content.
The continuum model described so far applies to bicrystal configurations with one GB ($\Sigma$). However, it can be considered for microstructures featuring a network of grain boundaries by defining parameters and geometrical properties accordingly. This is conveniently realized in the implicit description of GBs below, allowing us also to cope easily with complex morphological evolution (typical of grain growth).
@@
$\dot{\eta}_{i}=\sum_{j}^{N}M_{ij}\left(\frac{\delta F}{\delta\eta_{j}}-\frac{\delta F}{\delta\eta_{i}}+f_{\rm elastic}\right),$ (S-5)
@@
$\frac{\delta F}{\delta\eta_{j}}-\frac{\delta F}{\delta\eta_{i}}=f_{\rm curvature}+f_{\rm chemical},$ (S-6)
$f_{\rm curvature}=\Gamma_{ij}\left(\eta_{j}\nabla^{2}\eta_{i}-\eta_{i}\nabla^{2}\eta_{j}-\frac{\pi^{2}}{2\epsilon^{2}}(\eta_{j}-\eta_{i})\right),$ (S-7)
$f_{\rm chemical}=\frac{\pi}{\epsilon}\sqrt{\eta_{i}\eta_{j}}(\psi_{j}-\psi_{i}),$ (S-8)
$f_{\rm elastic}=\boldsymbol{\tau}\cdot\boldsymbol{\beta}_{ij}|\nabla\eta_{j}|(\hat{\mathbf{n}}(\eta_{i})\cdot\hat{\mathbf{n}}(\eta_{j})),$ (S-9)
@@
Numerical simulations are performed exploiting a finite difference spatial discretization of (S-5) and semi-implicit time integration. In practice, the domain is discretized into rectangles with edges $\mathrm{d}x_{1}$ and $\mathrm{d}x_{2}$. In all current PF simulations, $\mathrm{d}x_{1}=\mathrm{d}x_{2}=\epsilon/5$. The discretised gradient and Laplacian operators are
@@
where $l$ and $m$ are the index of the element along the $\mathbf{e}_{1}$ and $\mathbf{e}_{2}$ directions. All the simulations are performed under periodic boundary conditions in both $\mathbf{e}_{1}$ and $\mathbf{e}_{2}$ directions.
@@
We generate 30 samples for the PF simulations with different initial configurations. The initial configuration for grain growth simulations consists of 1000 circle nuclei with different orientations within a background matrix. A chemical potential difference between the nuclei and the matrix is applied so that the nuclei can rapidly consume the latter and densify into a polycrystal microstructure. We assume that the 2D polycrystals are [110]-textured in face-centered cubic metals; i.e., two sets of closely-packed and low-energy GB planes exist and are perpendicular to each other in the CSL lattices of [110]-tilt GBs. If we set the orientation $[001]\,||\,\mathbf{e}_{x}$, $[1\bar{1}2]\,||\,\mathbf{e}_{y}$ and $[110]\,||\,\mathbf{e}_{z}$ as the reference, according to crystal symmetry, the orientation angles of all grains should be in range of [0, $\pi$).
For the cases of isotropic mean curvature flow and those considering GB shear coupling (shown in Figs. 1-3 and 4f-h in main text), the GB energy ($\gamma$) and mobility ($M$) of all GBs are set as $\gamma=1$ and $M=4\pi^{2}/8\epsilon\approx 1$, where $\epsilon=5\mathrm{d}x$ is the interface thickness in the PF framework. The shear coupling factor of the reference interfaces of a certain arbitrarily-curved GB is determined by their misorientation angle ($\theta$) according to [4]:
@@
We note that shear coupling factors measured in polycrystals tend to be significantly smaller than those found in bicrystals [9, 10, 11, 4]. We account for this difference by introducing a shear coupling scaling factor, $S=1/30$. While the value of $S$ is chosen arbitrarily, it is of the correct order of magnitude compared with experimental observations. Nonetheless, the shear coupling factor expression, Eq. (S-12), retains the dependence of $\beta$ on bicrystallography consistent with theory, MD simulations, and experiments. The source of the smaller shear coupling in polycrystals compared with bicrystals is associated with the constraints imposed by other grains (i.e., triple junctions) in a polycrystal that activate additional disconnection modes [12, 13, 14].
@@
As for the [001]-textured idealized microstructure in copper, the initial configuration is also composed of square grains A and B and octagon grains C and D, with grains A and B sharing the same orientation, i.e., $[100]\,||\,\mathbf{e}_{x}$, $[010]\,||\,\mathbf{e}_{y}$ and $[001]\,||\,\mathbf{e}_{z}$. Grain C has a misorientation angle of -22.6 away from grains A and B. Grain D has a misorientation angle of 28.1 away from grains A and B. In this sense, the GBs between grains A and C, B and C are $\Sigma 13[100]$ asymmetric tilt GBs, whilst the GBs between grains A and D, B and D are $\Sigma 17[100]$ asymmetric tilt GBs. MD simulations are conducted with the EAM copper interatomic potential [15] and the same simulation methodology mentioned in [12]. In the PF simulation, the shear coupling factors are also assigned to the individual GBs based on their misorientations  [16, 11].
@@
In Figs. 4h,i of the main text, we introduce the misorientation-dependent GB energy. The formulation of GB energy follows the generalized GB energy functional proposed by Bulatov, Reed, and Kumar [17].
@@
Figure S-1: Impact of anisotropic mobility on the correlation between velocity and GB mean curvature: (a) Misorientation-dependent GB mobility from Eq. (S-15) and (c) (a) Randomly perturbed GB mobility resembling that reported in Ref. [18]. (b, d) Scatter plots (small points) of GB velocity vs. mean curvature for individual GBs obtained in grain growth considering anisotropic GB mobility in (a, c), respectively. The color of each small point represents the local density of measurements, with redder points indicating a higher concentration of grain boundaries (GBs) exhibiting similar velocities and curvatures. The open circles represent GB velocities averaged over GB segments with (approximately) the same mean curvature. Note that (d) is the same as Fig. 4(l) in the main text.
Figure S-2: Results on anisotropic grain growth considering anisotropic GB energy, mobility and triple junction drag effect: (a) Mean grain size $\langle R^{2}\rangle$ vs. time; (b) Mean grain size $\langle R\rangle$ vs. time. Scatter plots of grain area change $\delta R^{2}/\delta t$ vs. number of neighbours $N$ of grains during grain growth considering: (c) anisotropic GB energy; (d) additional smaller triple junction mobility; (e) anisotropic GB mobility; and (f) perturbed anisotropic GB mobility.
Figure S-3: Illustration of the GB replacement happening in one typical grain. The GBs are colored based on their GB energy.
The triple junction (TJ) drag effect is incorporated by assigning a lower phase-field (PF) kinetic coefficient in the TJ region (i.e., lower TJ mobility) compared to the grain boundary (GB) region (i.e., GB mobility) [19]. If a grid point is located inside the grain $i$ or inside the grain $j$ or at a GB between them, the condition $\eta_{i}(l,m)+\eta_{j}(l,m)=1$ holds. At TJs, instead, $\eta_{i}(l,m)+\eta_{j}(l,m)<1$. The mobility $M_{ij}$ at the corresponding grid point ($l,m$) in Eq. (S-5) is thus set to
$M_{ij}=M_{ij}^{\rm GB}\big{[}M_{ij}^{\rm TJ}+\exp(-500(\eta_{i}+\eta_{j}-1)^{2})(1-M_{ij}^{\rm TJ})\big{]}.$ (S-14)
If $\eta_{i}+\eta_{j}=1$ (at GBs), $M_{ij}=M_{ij}^{\rm GB}$ while if $\eta_{i}+\eta_{j}\ll 1$ (at TJs), $M_{ij}=M^{\rm GB}M_{ij}^{\rm TJ}$. In this work, we choose $M^{\rm TJ}=0.1$.
S-II.7 Anisotropic GB mobility
We consider two cases:
  1. (1)
    A misorientation- and inclination-dependent GB mobility reading
    $M(\theta)=\left\{\begin{aligned} &M_{0}\left[\frac{\theta}{\theta_{\rm c}}\left(1-\ln\frac{\theta}{\theta_{\rm c}}\right)\right],~{}~{}~{}~{}~{}\theta<\theta_{\rm c}\\ &M_{0},\quad\quad\quad\quad\quad\quad\quad\quad\quad~{}\theta\geq\theta_{\rm c}\end{aligned}\right.,$ (S-15)
    $M(\theta,\phi)=M(\theta)\cos^{2}\phi+M(\pi-\theta)\sin^{2}\phi,$
    where $\theta_{\rm c}$ is the critical misorientation angle and $M_{0}$ is the GB mobility of high-angle GBs (we choose $\theta_{\rm c}=\pi/6$ and $M_{0}=1$). The misorientation-dependent GB mobility is plotted in Fig. S-1a.
  2. (2)
    As shown in Ref. [18], misorientation-dependent GB mobility seems to be randomly dispersing around a constant value. Therefore, we here introduce a random perturbation to the GB mobility function with the values varying by one order of magnitude (with the similar characteristics to Figs. 4a and 6a in Ref. [18], see Fig. S-1c.
@@
In the main text, we mention that we conduct PF simulations of grain growth without GB shear coupling but with anisotropic GB energy, mobility, and triple junction (TJ) drag effect. Here, we provide more results on anisotropic grain growth in Fig. S-1 and Fig. S-2.
Figures S-1b,d provide the scatter plots of GB velocity vs. mean curvature for individual GBs obtained in grain growth considering anisotropic GB mobility. Figure S-1d reports the plot shown in Fig. 4h in the main text. In both cases, we show that anisotropic GB mobility can weaken the correlation between GB velocity and curvature, whilst its effect is still too weak to explain the loss of correlation observed in experiments and atomistic simulation fully illustrated in Fig. 4 in the main text.
Figures S-2a and b give the grain growth law of anisotropic grain growth with and without the TJ drag effect. The parabolic grain growth law is generally still obtained in the case of anisotropic grain growth. With additional TJ drag (i.e., smaller TJ mobility, see S-II.6), the grain growth exponent $n$ ($R^{n}-R_{0}^{n}=kt$) reduces from 2 (i.e., parabolic law) to 1.32. When the grain growth exponent $n=1$, the grain growth is controlled by TJ motion. Therefore, with the current TJ mobility (one order of magnitude smaller than the GB mobility), TJ effect is already clear.
@@
Figures S-2c-f show the scatter plots of grain size changes vs. the number of neighbours. Although the grain size changes vs. the number of neighbours relation has a larger dispersion than that obtained in isotropic mean curvature flow, the data distributions (and the shapes of the upper and lower limit) are still largely different from those in experiments. However, clear linear distributions are found, and the only difference is the reduction of the corresponding slope.
@@
Previous studies [20, 19] attributed the lack of correlation between grain boundary (GB) velocity and mean curvature in polycrystalline Ni and $\alpha$-Fe to GB replacement driven by GB energy anisotropy. However, we demonstrate in this work that GB velocity remains correlated with mean curvature, even when accounting for GB energy anisotropy and including the GB replacement mechanism as illustrated in Fig. S-3. Therein, high-energy boundaries 1 and 3 disappear during evolution, while the low-energy boundaries 2 and 7 increase in length. The grain size increases less than 1% during this procedure. This phenomenon can be described as ”GB replacement” according to the definition in [20, 21].